Show that lim π‘›β†’βˆž 1 /𝑛 [sin( πœ‹/ 𝑛 ) + sin( 2πœ‹/ 𝑛 ) + … + sin( π‘›πœ‹/ 𝑛 )] = 2/ πœ‹ .

1
Expert's answer
2021-10-15T09:48:00-0400

Solution:

"\begin{aligned}

&\lim _{n \rightarrow \infty} \frac{1}{n}\left[\sin \frac{\pi}{n}+\sin \frac{2 \pi}{n}+\ldots+\sin \frac{(n-1) \pi}{n}\right] \\

&\quad=\lim _{n \rightarrow \infty}\left\{\frac{1}{n} \sum_{r=1}^{n-1} \sin \left(\frac{r \pi}{n}\right)\right\}= \int_{0}^{1} \sin \pi x d x \\

&\quad=\left[\frac{-\cos \pi x}{\pi}\right]_{0}^{1}

\end{aligned}"

"=\dfrac1{\pi}[-\cos \pi+\cos 0]
\\=\dfrac1{\pi}[1+1]
\\=\dfrac2{\pi}"

Hence, proved.